### v1

# Map2Text: New Content Generation from Low-Dimensional Visualizations

Abstract.
Low-dimensional visualizations, or ”projection maps” of datasets are widely used across scientific research and creative industries as effective tools for interpreting large-scale and complex information. These visualizations not only support understanding existing knowledge spaces but are often used implicitly to guide exploration into unknown areas. While powerful methods like TSNE or UMAP can create such visual maps, there is currently no systematic way to leverage them for generating new content. To bridge this gap, we introduce Map2Text, a novel task that translates spatial coordinates within low-dimensional visualizations into new, coherent, and accurately aligned textual content. This allows users to explore and navigate undiscovered information embedded in these spatial layouts interactively and intuitively. To evaluate the performance of Map2Text methods, we propose Atometric, an evaluation metric that provides a granular assessment of logical coherence and alignment of the atomic statements in the generated texts. Experiments conducted across various datasets demonstrate the versatility of Map2Text in generating scientific research hypotheses, crafting synthetic personas, and devising strategies for testing large language models. Our findings highlight the potential of Map2Text to unlock new pathways for interacting with and navigating large-scale textual datasets, offering a novel framework for spatially guided content generation and discovery.
Textual Visualization, Spatially Guided Content Generation, Text Generation Evaluation.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
[…]
Illustration of Map2Text task on crafting synthetic persona. Refer to caption
Figure 1. Illustration of a Map2Text task which generates new content based on a 2D projection map. The figure provides an example of generating new red-teaming strategies for testing large language models. Given a low-dimensional visualization precomputed from certain text corpora and a query position (green dot), Map2Text challenges a model to generate new, coherent, and meaningful textual content that would correspond to that location in the visualization.
1. Introduction
Low-dimensional visualizations, or “projection” maps, are powerful tools that help users understand large textual corpora, much like maps aid in navigating physical landscapes. Typically, these visualizations are created by first generating high-dimensional embeddings using text embedding models (e.g.,  (Muennighoff et al., 2022)), followed by dimension reduction algorithms such as TSNE (Van der Maaten and Hinton, 2008), LargeVis (Tang et al., 2016), or UMAP (McInnes et al., 2018). This process results in a 2D (or 3D) map that preserves semantic relationships between texts, enabling users to intuitively explore complex datasets and identify interesting patterns (Low et al., 2020; Noichl, 2021; Yang et al., 2022; Tschisgale et al., 2023; Wang et al., 2022). This utility raises an intriguing question: much like how people use geographical maps to explore new territories, can we use these visualizations not only to navigate existing knowledge but also to generate new content?
This is the central idea behind our proposed task, Map2Text. More specifically, Map2Text aims to generate textual content corresponding to a given position within a low-dimensional visualization map. The fundamental assumption that Map2Text relies on is that semantically similar texts are positioned close together on a low-dimensional visualization, creating a “map” that conveys rich contextual information beyond the individual texts themselves. By leveraging this spatial context, Map2Text aims to guide the generation of new, coherent, and meaningful content.
Many downstream applications could benefit from Map2Text. In scientific idea generation (Gu and Krenn, 2024; Zhou et al., 2024; Wang et al., 2023), researchers can select positions on the visualization map where concepts are converging or underexplored. By generating text corresponding to these research gaps, Map2Text can inspire novel research directions and foster innovation. Another promising application is to test the safety and robustness of systems, including large language models (LLMs), which continuously benefits from new adversarial tests (Ganguli et al., 2022; Perez et al., 2022). Typically, developing new red teaming strategies relies on expert knowledge, which requires intensive learning processes and is not scalable. With Map2Text, security experts could navigate visualizations of existing adversarial attempts and generate new strategies at scale to identify and mitigate vulnerabilities in LLMs. Overall, Map2Text could be a valuable tool for any field that relies on continuous innovation and new content generation, through enabling the exploration into uncovered regions in an information space.
The Map2Text task presents several significant challenges. First, the process of reducing high-dimensional text embeddings to a two-dimensional map often introduce distortions and ambiguities in spatial representation. These artifacts may lead to inaccuracies when mapping back from a 2D point to the high-dimensional semantic space, affecting the relevance and coherence of the generated text (Rudin et al., 2022). Second, modeling text generation conditional on specific locations in the embedding space is non-trivial. Developing models that effectively integrate spatial information into the text generation process is essential to ensure that the output aligns with the intended semantics (Hu et al., 2017). Third, evaluating the quality and relevance of the generated content poses a challenge. Traditional evaluation metrics may not adequately capture the alignment between the generated text and its intended position on the map, necessitating the development of new evaluation frameworks (Goyal et al., 2022). Fourth, scalability and computational resource constraints are significant considerations. Handling large textual corpora and generating real-time responses based on user-selected points on the map can be computationally intensive, requiring optimization of algorithms for efficiency without compromising accuracy. Addressing these challenges is critical for the successful implementation and practical application of the Map2Text task.
Along this direction, we make the following initial contributions:
  • We introduce Map2Text, a novel task of generating text given positions on a visualization map, transforming existing visualization maps into tools for new content creation [^1] Code to replicate our experiments can be found at the following anonymous repo: https://anonymous.4open.science/r/map2text-27FD.
  • We propose Atometric, a reasoning-based evaluation metric that uses atomic statements to evaluate text generation quality at multiple levels of strictness [^2] https://anonymous.4open.science/r/atometric-17A3/.
  • We demonstrate Map2Text’s potential on four datasets, benchmarking the performance of seven candidate models and human baselines. We obtain interesting findings and showcase the potential of Map2Text across domains.
Organization
The remaining sections are organized as follows: Section 2 defines the problem formulation of Map2Text. Section 3 introduces the proposed evaluation metric, Atometric . Section 4 presents the experimental setup and results across multiple datasets. Section 5 reviews related work, and finally, Section 6 discusses some limitations of the paper and outlines future research directions.
Figure 2. An illustrative example of Atometric on generating new red teaming strategies for LLMs. Atometric decomposes the generated text into a set of atomic statements, which are individually compared against the reference text to evaluate how well each statement is supported under different levels of strictness—thereby measuring the “precision” of the generated text. Conversely, the reference text can be decomposed and compared against the generated text to assess “recall”. Both decomposition and verification processes are automated using LLMs with a prompt template, detailed in Appendix B.
2. Map2Text: A New Task
In this section, we introduce the formulation of Map2Text and its evaluation framework.
2.1. Problem Definition
Let ${\mathcal{V}}$ represent a low-dimensional projection of a text corpus $\mathcal{D}=\{s_{1},s_{2},\ldots,s_{N}\}$, where $s_{i}\in{\mathcal{S}}$ is the $i$-th text entry and ${\mathcal{S}}$ denotes the set of all possible text strings. For each text entry $s_{i}$, ${\mathcal{V}}$ includes a corresponding low-dimensional (e.g., 2D) position ${\bm{x}}_{i}\in{\mathbb{R}}^{2}$, forming a set of positions $\mathcal{X}=\{{\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{N}\}$. The visualization ${\mathcal{V}}$ is typically derived from $\mathcal{D}$ using a mapping function $f:{\mathcal{S}}\to{\mathbb{R}}^{2}$ that jointly maps each text entry $s_{i}$ to its position ${\bm{x}}_{i}$. This mapping function $f$ typically integrates a high-dimensional text embedding model and a dimensionality reduction algorithm. However, in many practical situations, the visualizations are precomputed and the details of the mapping function $f$ are unavailable to the users. Even when $f$ is known, reconstructing the visualization pipeline from scratch can be computationally expensive and subject to randomness. Thus, we assume no access to $f$ but rely on the assumption of locality preservation—points close to each other in the low-dimensional space $\mathcal{X}$ correspond to semantically similar text in the high-dimensional space ${\mathcal{S}}$. This allows us to use proximity in the 2D projection as a guide for generating contextually relevant text without requiring the original mapping function $f$.
Definition
The Map2Text task is defined as follows: given a query position ${\bm{x}}_{q}\in\mathbb{R}^{2}$ that has not been occupied by an existing text entry, a model $g$ is tasked with generating a text entry $\hat{s}$ that corresponds to ${\bm{x}}_{q}$. Specifically, the generated text $\hat{s}$ should be semantically and logically aligned with the ground truth text $s\in{\mathcal{S}}$. Here, a ground truth $s$ is a text entry that would have been mapped to ${\bm{x}}_{q}$ by the original mapping function $f$. Thus, the objective of the Map2Text task is to effectively learn an inverse mapping $g=\widehat{f^{-1}}:{\mathbb{R}}^{2}\to{\mathcal{S}}$ that reconstructs the relationship between the provided visualization and text entries.
2.2. Potential Solutions
Given the typical pipeline of $f$, there are multiple natural designs, each offering a distinct process on how to reconstruct text entries from low dimensional space.
The most straightforward approach is to directly learn an inverse mapping from the 2D positions to the corresponding text entries, requiring the model to memorize the entire visualization map, associate each position with its corresponding text, and generalize the learned relationship to unoccupied positions. This could be achieved by fine-tuning a pre-trained language model that takes the input of a 2D coordinate and outputs a text string. This process can be described as $g:{\mathbb{R}}^{2}\to{\mathcal{S}}$.
If high-dimensional embeddings of the text corpus are available, we could introduce an intermediate step by mapping the 2D positions to a high-dimensional textual embedding space before reconstructing the text. This can be achieved using a neural network. The embedding is then decoded back into text using an embedding inversion model such as the vec2text model (Morris et al., 2023) (note that an embedding inversion model requires access to the original embedding model $f_{e}:{\mathcal{S}}\to{\mathbb{R}}^{H}$). This approach can be expressed as $g:{\mathbb{R}}^{2}\to{\mathbb{R}}^{H}\to{\mathcal{S}}$, where ${\mathbb{R}}^{H}$ denotes the $\underline{H}$igh-dimensional embedding space.
Another approach may first generate an initial textual description or of the 2D position. This description is then be used to generate the final text, which can be done through prompting a language model. In this case, the inverse mapping is decomposed into two stages: first, the model constructs a preliminary text prompt, then uses this prompt to generate the final output, which could be described as $g:{\mathbb{R}}^{2}\to{\mathcal{S}}\to{\mathcal{S}}$.
A more detailed description and analysis of these methods and their performance is provided in Section 4.
2.3. Evaluation Framework
Like other tasks of new content generation, evaluating the Map2Text task presents unique challenges due to the lack of ground-truth content at arbitrary, previously unexplored positions. While domain experts could compare the quality and accuracy of generated contents given specific query positions, this online process is not scalable and thus unsuitable for automatic benchmarking.
To overcome this, we adopt an offline benchmarking approach by partitioning a precomputed visualization ${\mathcal{V}}$ into a training set and a testing set. In this approach, the training set includes both original text entries and their corresponding positions in the 2D map. This allows the model to understand the relationship between the text and positions. In contrast, for the testing set, only the positions are made available to the model, serving as query points for which the model generates the associated text. The generated text is then compared against the ground-truth text held out from the original visualization, which serves as a gold standard reference.
This framework addresses the scalability issue of human judgment but faces another challenge in aligning the generated content with the reference text. For instance, the generated text might use different wording or phrasing while still conveying the same underlying idea as in the reference. Traditional lexical-level metrics may fail to capture such equivalence, penalizing otherwise valid outputs. This motivates the proposal of Atometric, an evaluation metric designed to assess semantic and logical alignment between the generated and reference texts across varying levels of strictness.
3. Atometric: A New Evaluation Metric
A key challenge of evaluating Map2Text is to identify an appropriate metric that accurately assesses the alignment between the generated content $\hat{s}$ and the gold standard reference $s$ in our specific context. We outline the essential requirements for such a metric and introduce a novel metric tailored to meet these criteria.
3.1. Motivation
Unlike traditional content generation tasks, which aims to replicate ground-truth texts, Map2Text requires evaluating whether the generated content is capable of inspiring the user to explore the map in a similar way—despite how much the surface expressions match or differ, the generated content must accurately reflect the semantic relationships and logical structures of reference content.
Existing evaluation metrics struggle to capture the semantic and logical complexity required for the Map2Text task. Traditional metrics such as ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002) rely on lexical-level n-gram overlaps, penalizing legitimate variations in phrasing and structure even when the generated content accurately conveys the same meaning. Recent metrics like BERTScore (Zhang et al., 2019) and BLEURT (Sellam et al., 2020) address some of these limitations by using contextual embeddings to measure semantic similarity. However, these metrics lack interpretability, as they offer a high-level similarity scores without explaining the nuances of why certain outputs are deemed similar or dissimilar. Moreover, these metrics evaluate holistic sentence- or paragraph-level similarity, neglecting the internal coherence of fine-grained statements. This limitation is crucial for Map2Text , where maintaining the precise relationships and logical structures between entities is critical.
3.2. Features
To address the limitations of existing metrics, we propose Atometric, a novel metric specifically designed to evaluate the semantic and logical coherence of generated content in the Map2Text task. An illustrative example of Atometric is shown in Figure 2. Atometric has the following features:
Atomic Statement Alignment.
Inspired by FActScore (Min et al., 2023), Atometric focuses on the alignment of atomic statements—the fundamental units of meaning that express relationships between entities and actions within a text. By directly evaluating these atomic statements instead of the complex full texts or fragmented ngrams, Atometric assesses whether the generated content preserves the logical relationships and semantic properties present in the reference.
Multiple Levels of Strictness.
Atometric introduces multiple strictness levels to capture nuanced levels of alignment. Inspired by definitions in textual entailment (Bowman et al., 2015a; Sadat and Caragea, 2022; Williams et al., 2018), these strictness levels offer varying degrees of relevance, from direct support to more relaxed criteria such as topic relevance. This flexibility ensures that the metric can accommodate the inherent variability of natural language while remaining faithful to the core semantic content.
Comprehensive Evaluation Metrics:
Atometric provides precision, recall, and F1-score measurements that jointly evaluate the correctness and completeness of the generated content. Precision measures the logical consistency and accuracy of generated statements, while recall captures the extent to which the generated content covers all key statements of the reference. This approach ensures a thorough assessment of both semantic accuracy and content coverage.
3.3. Definition
The formal definition of Atometric is as follows. Given a pair of generated and reference text $(\hat{s},s)$, We first extracts a set of atomic statements from $\hat{s}$ and $s$ using an LLM, denoted as ${\mathcal{A}}_{\hat{s}}$ and ${\mathcal{A}}_{s}$, respectively. The two sets of atomic statements are further verified by another LLM at multiple levels of strictness $l$ to evaluate both precision and recall. Mathematically, Atometric precision, recall, and $F_{1}$-score at level $l$ is defined as:
(1) $\text{Atometric}_{P}$ $=\dfrac{1}{|{\mathcal{A}}_{\hat{s}}|}\sum_{a\in{\mathcal{A}}_{\hat{s}}}{\mathbb{I}}^{l}[a\,\text{is relevant to}\,s]$
(2) $\text{Atometric}_{R}$ $=\dfrac{1}{|{\mathcal{A}}_{s}|}\sum_{a\in{\mathcal{A}}_{s}}{\mathbb{I}}^{l}[a\,\text{is relevant to}\,\hat{s}]$
(3) $\text{Atometric}_{F1}$ $=\dfrac{2\times\text{Atometric}_{P}\times\text{Atometric}_{R}}{\text{Atometric}_{P}+\text{Atometric}_{R}}$
where ${\mathbb{I}}^{l}[\cdot]$ is an indicator function at strictness level $l$ that checks if an atomic statement is relevant to the corresponding reference or generated text in the sense of textual entailment (Sadat and Caragea, 2022; Bowman et al., 2015a; Williams et al., 2018). We show an overview of criteria in Table 1 and provide a detailed template in Appendix Figure 5 that could be easily adapted to different datasets with minimal change.
Level Criterion
Loose The reference and the atomic statement share related concepts or themes, and do not contradict each other.
Moderate The atomic statement can be logically inferred from the reference, or the atomic statement can support the reference.
Strict The reference explicitly states or clearly paraphrases the same information as the atomic statement.
Table 1. Overview of different strictness criteria. A bolded letter of each level is used as its abbreviation in later references. Details are provided in Appendix B.
Comment
Atometric builds on the strengths of existing evaluation metrics while addressing their limitations. Atometric Precision and Atometric Recall can be viewed as relaxations of BLEU and ROUGE, respectively, as they shift from strict n-gram matching to statement-level alignment. This allows generation with more flexible expression while preserving underlining semantic and logical relationships. Compared to BERTScore and BLEURT, which rely on embedding similarity, Atometric uses LLM-based multi-level relevance judgments, offering a more interpretable and adjustable evaluation. By focusing on atomic statements and incorporating varying strictness levels, Atometric provides a more nuanced and detailed assessment of content generation, particularly suited to tasks like Map2Text that require a balance between semantic accuracy and logical coherence.
4. Experiments
With the new evaluation metric, we are able to benchmark a set of simple yet effective candidate methods for Map2Text, instantiating the outlined solutions in Section 2.2, and compare them with a human baseline and a strong dummy baseline.
4.1. Candidate Methods
Fine-Tuning (FT)
The most straightforward approach is to directly learn the inverse mapping $g:{\mathbb{R}}^{2}\to{\mathcal{S}}$ by fine-tuning a pre-trained model. We fine-tune the 70B Llama 3.1 model using LoRA (Dubey et al., 2024), where the model takes a 2D coordinate as input and outputs the corresponding text. For instance, an input query would be, ”Convert the coordinate to text: [9.1054, 10.1339].”
Embedding Inversion
If high-dimensional embeddings are available, we can decompose $g$ into two separate mappings: $g_{1}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{H}$ and $g_{2}:{\mathbb{R}}^{H}\to{\mathcal{S}}$. For the first mapping, $g_{1}$, we use a k-NN interpolation algorithm to project the 2D coordinates into the $H$-dimensional embedding space. While more complex methods could be employed for $g_{1}$ (e.g., training a neural network), we select this straightforward approach as it requires no additional training and demonstrates comparable empirical performance. The second mapping, $g_{2}$, then converts the high-dimensional embeddings into text using the vec2text model (Morris et al., 2023). Since we use the pre-trained vec2text model without any additional fine-tuning, it is limited to supporting embeddings generated specifically by text-embedding-ada-002.
Retrieval-Augmented Generation (RAG)
Prompting LLMs has shown strong performance across various tasks (Lewis et al., 2020). However, a straightforward zero-shot prompting approach—like the query mentioned in the fine-tuning example above—is not feasible as a pre-trained LLM lacks inherent knowledge of the visualization map. To address this limitation, we incorporate a set of Retrieval-Augmented Generation (RAG) methods on GPT-4o to provide relevant context and enhance the model’s understanding of the map. We use the same notation in Section 2 to illustrate each method. For each query position ${\bm{x}}_{q}\in{\mathbb{R}}^{2}$,
  1. (1)
    (1st-order RAG) We retrieve the texts corresponding to its nearest neighbors $\mathcal{N}(\mathbf{x}_{q})\subset\{1,\dots,N\}$ and use this information to prompt the LLM to generate new content $\hat{s}$.
  2. (2)
    (1st-order RAG with few-shot examples) Building on (1), we add few-shot examples. Each few-shot example consists of texts from $\mathcal{N}(\mathbf{x}_{i})$ and the corresponding $s_{i}$ for some $i\in\mathcal{I}$, where $\mathcal{I}$ is a pre-selected set.
  3. (3)
    (2nd-order RAG with few-shot examples) Building on (2), we replace the pre-selected set $\mathcal{I}$ with $\mathcal{N}(\mathbf{x}_{q})$. Intuitively, this means we consider information from ”2-hop” neighbors of the query position.
  4. (4)
    (1st-order RAG with chain-of-thought examples) Building on (2), we add additional chain-of-thought reasoning (Wei et al., 2022) as auxiliary information.
Implementation of each method can be found in Appendix D.
4.2. Datasets
We evaluate the candidate models on four diverse datasets to demonstrate the broad application of Map2Text.
Dataset Name # Entries Avg. Len. Vis. Pipeline
Persona 100k 27.10 gte-v1.5 + UMAP
Red-Teaming Strategies 39k 84.06 ada-002 + UMAP
CS Research Idea 149k 41.57 ada-002 + LargeVis
CS Research Context 150k 40.96 ada-002 + LargeVis
Table 2. Basic information of datasets and visualizations. Average length is calculated using tiktoken. A bolded letter of each dataset is used as its abbreviation in later references.
Persona
We use an existing visualization [^3] https://huggingface.co/datasets/argilla/FinePersonas-v0.1-clustering-100k of a subset of the PersonHub dataset (Chan et al., 2024) where each text entry provides a brief description of a synthetic persona. Map2Text can be applied to generate new persona descriptions based on different regions of the map, enabling the creation of diverse character profiles.
Red-Teaming Strategies
We construct a summarization dataset [^4] Details of the construction are provided in Appendix C.2. from the LLM Red Teaming dataset by Ganguli et al. (2022) which comprises human-generated and annotated red-teaming dialogues aimed at identifying vulnerabilities in LLMs. Given a 2D visulization of these summaries, Map2Text can help generate unexplored red-teaming strategies, enhancing the robustness testing of LLMs.
CS Research Ideas & Context
We create 2D maps of the MASSW dataset (Zhang et al., 2024), where each entry corresponds to the research context or the key idea of a computer science publication. A good Map2Text model can assist scientific innovation by generating new research ideas or new context of emerging topics from unexplored regions on the map, helping researchers to identify gaps in current research and inspire novel directions for scientific inquiry.
The statistics of these datasets are summarized into Table 2. Note that the visualization pipelines for these datasets are different and are hidden from candidate Map2Text methods (except for embedding inversion), and a robust model should be capable of generalizing across different configurations.
4.3. Experiment Setup
Data Split
Each dataset is divided into training and testing sets, as described in Section 2.3, with 200 samples consistently held out for testing across all datasets. We apply a random split for the Persona and Red-Teaming Strategies datasets and a time-based split for the CS Research Ideas & Context datasets, where the most recent 200 entries (all in Dec 2023) are held out for testing, simulating the scenario of real-world research exploration.
Evaluation Metrics
We prioritize our proposed evaluation metric, Atometric , as it is specifically designed to capture the nuanced semantic and contextual alignment. We report Atometric-F1 across three levels of strictness: loose (L), moderate (M), and strict (S), as well as Atometric-Precision (M) and Atometric-Recall (R) for a more comprehensive picture [^5] We use gpt-4o-2024-05-13 as the backbone LLM in Atometric.. In addition to Atometric, we include traditional evaluation metrics—BERTScore (F1) (Zhang et al., 2019), BLEURT (Sellam et al., 2020), METEOR (Lavie and Agarwal, 2007), and ROUGE-2 (Lin, 2004)—as reference points for comparison. Evaluation on some extra metrics such as BLEU and other Atometric scores are provided in Appendix A.1.
4.4. Baselines
We begin by analyzing the evaluation results of a human baseline and a simple yet effective dummy baseline, providing a foundation for comparing the performance of the candidate AI methods.
Human Baseline
We recruit human annotators to conduct the Map2Text task using the Persona dataset, as this dataset does not require special domain expertise. Detailed settings for this study are provided in Appendix A.2. In brief, human annotators are given access to a partial visualization containing the query location and they are allowed to view the textual content of all visible data points. No constraints are imposed regarding their generation strategy, allowing them complete flexibility in navigating the visualization and generating content based on the map. We expect the human baseline to represent a strong benchmark compared to AI methods.
EchoNearest Baseline
For each query ${\bm{x}}_{q}$, the EchoNearest model simply retrieves and outputs the content of its closest neighbor in the 2D space. This simple baseline operates under the assumption that semantically similar content is clustered closely in the visualization map, making the nearest neighbor a relevant choice for responding to the query, though it lacks novelty. This approach mimics a “lazy” human strategy for content generation which could be effective when novelty is not required. As our gold standard references all exhibit some level of novelty compared to their neighbors, we expect it to perform below both the human baseline and more advanced AI methods when evaluated using proper metrics.
Analysis
Metric Human EchoNearest
P P R I C
Atometric-F1 (L) 0.901 0.884 0.783 0.430 0.538
Atometric-F1 (M) 0.605 0.556 0.586 0.191 0.105
Atometric-F1 (S) 0.302 0.268 0.282 0.098 0.045
Atometric-P (M) 0.610 0.548 0.598 0.199 0.103
Atometric-R (M) 0.601 0.565 0.575 0.184 0.107
BERTScore F1 0.894 0.898 0.902 0.857 0.862
BLEURT Scores 0.431 0.454 0.436 0.325 0.323
METEOR 0.232 0.306 0.428 0.160 0.170
ROUGE-2 0.110 0.117 0.231 0.032 0.034
Table 3. Evaluation results for the human and EchoNearest baselines. (L), (M), and (S) denote “loose”, “moderate”, and “strict” for the strictness level of Atometric.
The evaluation results for the human and EchoNearest baselines are provided in Table 3. Several observations can be made.
First, under all Atometrics, the human baseline outperforms the EchoNearest model, as expected. However, under conventional metrics, the human baseline scores lower than the EchoNearest model. This discrepancy likely arises because conventional metrics emphasize lexical-level text similarity, favoring the EchoNearest model’s repetition of existing entries. In contrast, human responses, while more contextually relevant, tend to vary in language style and phrasing, which conventional metrics fail to capture. This validates the strength of Atometric in recognizing the semantic and contextual richness of human-generated content.
Second, the four datasets exhibit varying levels of complexity. The EchoNearest model performs relatively well on the Persona and Red-Teaming Strategies datasets, suggesting that the text entries clustered closely in the visualization are highly similar, making these tasks easier to address. In contrast, the model’s performance on the CS Research Idea and CS Research Context datasets is notably lower. This indicates that these tasks require generating more complex or novel content, which demands a deeper understanding of the research domain, beyond simple proximity in the 2D space.
In addition, the Atometric results across different levels of strictness reveal subtle distinctions between research contexts and research ideas, showcasing the strength of Atometric in capturing varying levels of sensitivity. At lower strictness levels, Atometric F1 (L) scores are higher for the CS Research Contexts. While each paper addresses a unique problem, the overarching topics are often related, leading to greater alignment in topic-level relevance. In contrast, at medium and high strictness levels (M and S), research ideas exhibit higher Atometric scores. This occurs because research ideas, while varied, frequently share concrete methodologies or designs, which stricter criteria capture more effectively. By distinguishing these nuanced differences, Atometric demonstrates its ability to assess content sensitivity at multiple levels.
4.5. Benchmark Results
Dataset Metric GPT-4o Llama 3.1 Embedding
CoT-RAG(1) FS-RAG(1) RAG(2) RAG(1) FT RAG(1) Inversion
Persona Atometric-F1 (L) 0.893 0.852 0.824 0.822 0.822 0.783 -
Atometric-F1 (M) 0.602 0.494 0.464 0.454 0.463 0.390 -
Atometric-F1 (S) 0.254 0.183 0.168 0.158 0.194 0.117 -
Atometric-P (M) 0.582 0.460 0.449 0.430 0.459 0.365 -
Atometric-R (M) 0.623 0.533 0.480 0.481 0.467 0.420 -
BERTScore F1 0.898 0.895 0.893 0.892 0.893 0.887 -
BLEURT Scores 0.474 0.454 0.441 0.433 0.406 0.416 -
METEOR 0.329 0.290 0.263 0.260 0.251 0.268 -
ROUGE-2 0.116 0.096 0.078 0.086 0.098 0.074 -
[…]
Atometric-F1 (M) 0.514 0.480 0.451 0.414 0.497 0.437 0.428
Atometric-F1 (S) 0.202 0.200 0.168 0.144 0.185 0.167 0.126
Atometric-P (M) 0.479 0.445 0.397 0.358 0.508 0.384 0.329
Atometric-R (M) 0.555 0.521 0.520 0.491 0.486 0.507 0.612
BERTScore F1 0.904 0.902 0.901 0.897 0.894 0.898 0.877
BLEURT Scores 0.446 0.439 0.435 0.417 0.407 0.430 0.394
METEOR 0.433 0.424 0.412 0.371 0.390 0.405 0.312
ROUGE-2 0.250 0.250 0.244 0.232 0.203 0.235 0.114
[…]
Atometric-F1 (M) 0.213 0.220 0.214 0.202 0.163 0.210 0.180
Atometric-F1 (S) 0.091 0.098 0.102 0.099 0.084 0.100 0.096
Atometric-P (M) 0.222 0.240 0.272 0.250 0.203 0.263 0.193
Atometric-R (M) 0.205 0.202 0.176 0.169 0.136 0.175 0.169
BERTScore F1 0.860 0.860 0.864 0.864 0.858 0.863 0.861
BLEURT Scores 0.367 0.365 0.350 0.348 0.302 0.337 0.334
METEOR 0.198 0.194 0.153 0.142 0.153 0.159 0.173
ROUGE-2 0.041 0.037 0.034 0.033 0.029 0.034 0.037
[…]
Atometric-F1 (M) 0.144 0.144 0.145 0.129 0.094 0.110 0.133
Atometric-F1 (S) 0.055 0.057 0.055 0.054 0.032 0.039 0.055
Atometric-P (M) 0.181 0.179 0.160 0.165 0.108 0.135 0.113
Atometric-R (M) 0.119 0.120 0.133 0.106 0.084 0.093 0.161
BERTScore F1 0.868 0.868 0.867 0.868 0.859 0.863 0.867
BLEURT Scores 0.340 0.345 0.341 0.339 0.286 0.323 0.340
METEOR 0.173 0.166 0.164 0.159 0.137 0.158 0.190
ROUGE-2 0.046 0.046 0.041 0.044 0.024 0.033 0.034
Table 4. Evaluation results of the candidate methods on the Map2Text task. The highest score(s) in each row are highlighted in bold and the second highest score(s) are in italic. Results of additional evaluation metrics (e.g., Atometric Precisions and Recalls) are provided in Appendix A.1. (L), (M), and (S) denote “loose”, “moderate”, and “strict” for strictness level. Note that the embedding inversion method relies on vec2text (Morris et al., 2023), which only supports the text-embedding-ada-002 model. As a result, no scores are provided for the Persona dataset.
Table 4.5 reveals several patterns in the performance of the candidate AI methods in Section 4.1 across the different datasets. No single method consistently dominates across all metrics, reflecting the varying complexities and characteristics of the domains.
Analysis
For the Persona and Red-Teaming Strategies datasets, CoT-RAG(1) consistently outperforms other AI methods across most metrics. This suggests that incorporating more complex reasoning structures, such as chain-of-thoughts, enhances the ability of LLMs to better align with the query position. Additionally, fine-tuning (FT) on these datasets shows clear improvements on Llama 3.1 compared to applying RAG(1) on its untuned version. This indicates that fine-tuning enables the LLM to better capture the specific nuances of the relation between the map locations and text.
In contrast, for CS Research Ideas and CS Research Contexts, no single method clearly stands out. The best performer under most metrics is one of the GPT-4o based methods, indicating the importance of using a more capable foundation model. However, more complex prompting only leads to marginal improvements over the basic RAG(1) method. Interestingly, fine-tuning Llama 3.1 on these datasets appears to reduce its performance compared to the untuned RAG(1) model. One possible explanation is that these datasets contain highly specialized technical content, which may not align well with the general knowledge encoded in pre-trained LLMs. Without sufficient domain-specific data for fine-tuning, a model may lose some of its generalization capacity.
Additionally, the Embedding Inversion method demonstrates competitive performance, particularly under Atometric Recall, albeit with some loss in precision. On the Red-Teaming Strategies and CS Research Contexts, it achieves scores comparable to those of GPT-4o’s RAG(1) model, despite only employing a simple k-NN interpolation and a small pre-trained vec2text model (220M parameters) without additional training. This result highlights the potential of embedding inversion strategies for Map2Text tasks, when the originally embedding methods are accessible.
Surprisingly, no candidate models surpass the EchoNearest baseline on the Red-Team Strategies dataset. This is likely due to the concentrated distribution of the dataset, as shown in Appendix Figure 7—with examples being so close to one another that distinguishing them becomes difficult. Intuitively, when the map is sparser (i.e., KNNs are farther apart), targeting relevance is more challenging. Conversely, when the map is locally denser (i.e., KNNs are closer together), relevance is easier to achieve, but innovation becomes harder, an issue that could be addressed in future work.
Finally, the human baseline demonstrates superior performance compared to all candidate methods on the Persona dataset. Its primary advantage is the flexibility to navigate the visualization and extract high-level information, such as the distribution of topics across different areas and the directional relationships between nearest neighbors. This capability enables more accurate and contextually appropriate content generation, suggesting that there is still broad improvement space for better AI methods to incorporate similar high-level navigational strategies on the map.
5. Related Work
To the best of our knowledge, this is the first study on new content generation from low-dimensional projection maps. This new task is related to a few lines of research.
Low-Dimensional Visualization of Text Corpora
Low-dimensional data visualization is a versatile tool used across numerous fields to provide intuitive representations of large datasets (Sadiku et al., 2016). In public health, it assists in tracking disease trends and guiding data-driven decision-making (Kilimba et al., 2015). In renewable energy, visualization helps optimize energy production and consumption (Kumar and Goyal, 2016), while environmental science relies on geographic visualization to interpret ecological data (Bohman et al., 2015; MacEachren and Taylor, 2013). Fraud detection benefits from interactive visualizations that help identify suspicious patterns (Dilla and Raschke, 2015), and library management uses visualization techniques to manage resources and support strategic decisions (Murphy, 2013). A variety of methods are employed to achieve these visualizations. Textual embedding models (Muennighoff et al., 2022), such as OpenAI’s Ada 2 (OpenAI, nd) and the widely-used BERT (Devlin, 2018), transform text into dense vector representations that capture underlying semantic relationships. Dimensionality reduction algorithms like t-SNE (Van der Maaten and Hinton, 2008), LargeVis (Tang et al., 2016), and UMAP (McInnes et al., 2018) are commonly utilized to project the high-dimensional embeddings onto 2D or 3D maps, making complex patterns in the data space more accessible for analysis. While these visualizations primarily focus on interpreting the space of known knowledge, Map2Text extends the capabilities of the map by enabling the generation of new, relevant, and accurately aligned textual content. This shift bridges the gap between data interpretation and creative exploration, offering new possibilities for leveraging data visualizations to facilitate applications such as scientific discovery and innovation.
Controllable Text Generation
Controllable text generation (CTG) refers to the task of generating text according to the given controlled element  (Prabhumoye et al., 2020). It has been a critical focus in recent years and covers a wide range of tasks, such as Attribute-based Generation and Data Augmentation. In Attribute-based Generation, the goal is to generate text that aligns with specific attributes, such as topics, emotions, or keywords, allowing systems to produce interpretable and structured outputs tailored to user-specified needs. For instance, Khalifa et al. (2020) offers a method to manipulate attributes like topic by adjusting latent variables, improving control over the text’s structure and coherence. Similarly, Prabhumoye et al. (2020) examines a broader range of controllable attributes, demonstrating how various methods can enhance the generation process, leading to text that is more aligned with user goals. Data Augmentation is another important task within CTG. Works like Amin-Nejad et al. (2020) and Liu et al. (2020) have highlighted the effectiveness of CTG for data augmentation, where generated text supplements existing data by altering attributes or replacing specific entities, thus broadening the training set and improving model performance. While these works emphasize direct control over attributes such as topic or slot values, Map2Text shifts the focus to leveraging spatial relationship of text representations, using the contextual information in a visualization map to guide content creation.
Evaluation of Text Generation
The evaluation of text generation quality has traditionally relied on similarity-based metrics, broadly categorized into lexical- and semantic-based approaches. These methods have been widely adopted across tasks such as machine translation, summarization, and creative content generation. Lexical-based metrics focus on n-gram overlaps between generated and reference texts. BLEU emphasizes precision, making it a common choice for machine translation (Papineni et al., 2002), while ROUGE, which prioritizes recall, is frequently used for summarization tasks (Lin, 2004). METEOR (Lavie and Agarwal, 2007), another well-known metric, improves by incorporating synonyms and stemming to better handle linguistic variations. However, these metrics often fail to capture the deeper semantic meaning of text. To address these limitations, semantic-based metrics have been developed. Techniques such as BERTScore (Zhang et al., 2019) leverage contextual embeddings from BERT to capture the token-level similarity between sentences. BLEURT (Sellam et al., 2020), by fine-tuning models to predict human judgment, further aligns with human perception and measure beyond surface-level overlaps.
More recently, reasoning-based evaluations have emerged, utilizing the capabilities of LLMs to assess text coherence, consistency, and factual accuracy. Methods like GPTScore (Fu et al., 2023) employ LLM-predicted probabilities as quality scores, while G-Eval (Liu et al., 2023) engages LLMs in tasks such as Likert scale evaluations and pairwise comparisons. These approaches leverage the reasoning abilities of LLMs to evaluate content beyond traditional metrics. FActScore (Min et al., 2023) takes this a step further by breaking generated text into atomic facts and verifying each fact’s support against external knowledge sources. Building on these advancements, our proposed method, Atometric, offers a more comprehensive and interpretable evaluation by assessing both precision and recall of atomic statements at multiple levels of strictness. This approach captures the nuances of semantic relevance and logical coherence of the generated text, addressing gaps left by previous metrics.
6. Conclusion
In this work, we introduce Map2Text, a novel task for generating new textual content guided by specific positions within a 2D visualization map of a large text corpus. We propose a new evaluation metric, Atometric, to assess the fine-grained alignment of the generated text to gold standard references at varying levels of strictness. Through this metric, we demonstrate the effectiveness of multiple representative candidate methods for Map2Text on datasets in diverse domains. Our findings show the potential of Map2Text as a valuable tool for applications such as scientific idea generation, new persona generation, and LLM red teaming. This initial attempt calls for future research may explore more advanced methods and diverse applications of Map2Text.
6.1. Limitations and Future Work
We list several limitations and potential directions that could be addressed in future research:
Novelty of Generated Content:
Our current evaluation does not go beyond measuring similarity between the generated content and a given gold standard reference. For example, it does not consider the novelty of the output or the reference, which is crucial for applications such as research idea generation. One potential solution is to compute the similarity between the generated content and its nearest neighbors (where lower similarity indicates higher novelty) rather than the reference alone. In reality, the quality of the “gold” standard itself is often not guaranteed, as human-generated content can be highly relevant while still lack scientific validity or logical coherence. Future work may employ domain-specific criteria into the evaluation metrics or rely on specific downstream tasks.
Evaluation Limited to Known Positions:
Our evaluation framework is limited to positions on the map for which ground truth content exists but is held out. We do not evaluate the model’s ability to generate content at arbitrary unexplored positions on the map, which would unleash the full potential of Map2Text for open exploration and innovation into the information space. Without the gold stand reference, the quality of the generated new content has to be evaluated otherwise, through online human judgments or the utility of the generated content in downstream tasks, such as the effectiveness of a new red-teaming strategy in testing LLMs.
Optimizing Map2Text Methods:
There is considerable room of improvement beyond the candidate methods tested in this paper. An advanced Map2Text model may leverage global patterns or community structures in the map to guide the generation, beyond the local neighborhood. Another approach is to fine-tune LLMs to utilize spatial tokens that encode positional information directly within their embedding layers. This would allow an LLM to memorize the complex map structure and incorporate spatial context more effectively, while not harming their reasoning ability.
Appendix A Experiment Details
A.1. Results on Extra Metrics
We provide additional experiment results on a broader list of evaluation metrics as a complementary reference for Table 3 and 4.5. The results are shown in Table A.1, 6, and 7.
Dataset Metric GPT-4o Llama 3.1 Embedding
CoT-RAG(1) FS-RAG(1) RAG(2) RAG(1) FT RAG(1) Inversion
Persona Atometric Precision (loose) 0.908 0.853 0.822 0.815 0.822 0.758 -
Atometric Precision (moderate) 0.582 0.460 0.449 0.430 0.459 0.365 -
Atometric Precision (strict) 0.201 0.141 0.133 0.125 0.208 0.096 -
Atometric Recall (loose) 0.878 0.850 0.825 0.829 0.822 0.808 -
Atometric Recall (moderate) 0.623 0.533 0.480 0.481 0.467 0.420 -
Atometric Recall (strict) 0.342 0.261 0.229 0.214 0.181 0.148 -
BERTScore Precision 0.893 0.890 0.891 0.891 0.897 0.882 -
BERTScore Recall 0.903 0.900 0.895 0.894 0.890 0.892 -
BLEU (BLEU) 0.075 0.063 0.050 0.060 0.056 0.045 -
Cosine Similarity 0.885 0.879 0.871 0.869 0.868 0.857 -
ROUGE-1 0.360 0.334 0.311 0.309 0.315 0.298 -
ROUGE-L 0.281 0.262 0.244 0.248 0.258 0.231 -
ROUGE-Lsum 0.281 0.262 0.245 0.247 0.258 0.231 -
[…]
Atometric Precision (moderate) 0.479 0.445 0.397 0.358 0.508 0.384 0.329
Atometric Precision (strict) 0.182 0.180 0.147 0.142 0.188 0.156 0.093
Atometric Recall (loose) 0.744 0.718 0.724 0.695 0.688 0.715 0.826
Atometric Recall (moderate) 0.555 0.521 0.520 0.491 0.486 0.507 0.612
Atometric Recall (strict) 0.226 0.224 0.196 0.146 0.183 0.180 0.194
BERTScore Precision 0.904 0.903 0.903 0.901 0.894 0.899 0.879
BERTScore Recall 0.903 0.901 0.899 0.893 0.894 0.898 0.875
BLEU (BLEU) 0.231 0.230 0.221 0.198 0.184 0.214 0.076
Cosine Similarity 0.926 0.918 0.915 0.906 0.913 0.911 0.926
ROUGE-1 0.481 0.478 0.471 0.446 0.444 0.465 0.430
ROUGE-L 0.377 0.376 0.372 0.354 0.329 0.359 0.250
ROUGE-Lsum 0.378 0.376 0.372 0.354 0.330 0.360 0.250
[…]
Atometric Precision (moderate) 0.222 0.240 0.272 0.250 0.203 0.263 0.193
Atometric Precision (strict) 0.084 0.095 0.117 0.118 0.110 0.118 0.115
Atometric Recall (loose) 0.477 0.453 0.429 0.450 0.378 0.454 0.511
Atometric Recall (moderate) 0.205 0.202 0.176 0.169 0.136 0.175 0.169
Atometric Recall (strict) 0.099 0.102 0.090 0.086 0.068 0.086 0.082
BERTScore Precision 0.858 0.857 0.869 0.870 0.863 0.868 0.864
BERTScore Recall 0.862 0.863 0.859 0.858 0.853 0.858 0.858
BLEU (BLEU) 0.024 0.020 0.020 0.016 0.016 0.021 0.025
Cosine Similarity 0.836 0.836 0.833 0.834 0.822 0.832 0.839
ROUGE-1 0.221 0.225 0.205 0.205 0.206 0.214 0.227
ROUGE-L 0.162 0.160 0.157 0.158 0.153 0.156 0.168
ROUGE-Lsum 0.162 0.160 0.157 0.158 0.153 0.156 0.168
[…]
Atometric Precision (moderate) 0.181 0.179 0.160 0.165 0.108 0.135 0.113
Atometric Precision (strict) 0.060 0.070 0.062 0.077 0.035 0.045 0.049
Atometric Recall (loose) 0.620 0.588 0.580 0.569 0.478 0.537 0.682
Atometric Recall (moderate) 0.119 0.120 0.133 0.106 0.084 0.093 0.161
Atometric Recall (strict) 0.050 0.048 0.049 0.041 0.029 0.034 0.063
BERTScore Precision 0.871 0.873 0.872 0.874 0.864 0.869 0.870
BERTScore Recall 0.864 0.864 0.862 0.862 0.853 0.858 0.863
BLEU (BLEU) 0.028 0.030 0.028 0.028 0.017 0.021 0.029
Cosine Similarity 0.849 0.849 0.847 0.847 0.823 0.831 0.856
ROUGE-1 0.214 0.202 0.199 0.200 0.171 0.202 0.230
ROUGE-L 0.151 0.151 0.142 0.147 0.121 0.139 0.151
ROUGE-Lsum 0.151 0.151 0.142 0.147 0.121 0.140 0.151
Table 5. Additional evaluation results of the candidate methods on the Map2Text task. (L), (M), and (S) denote “loose”, “moderate”, and “strict” for strictness level.
Metric P R I C
Atometric Precision (loose) 0.882 0.802 0.453 0.542
Atometric Precision (moderate) 0.548 0.598 0.199 0.103
Atometric Precision (strict) 0.296 0.288 0.103 0.044
Atometric Recall (loose) 0.885 0.764 0.411 0.535
Atometric Recall (moderate) 0.565 0.575 0.184 0.107
Atometric Recall (strict) 0.267 0.279 0.093 0.045
BERTScore Precision 0.898 0.902 0.860 0.864
BERTScore Recall 0.898 0.902 0.855 0.860
BLEU (BLEU) 0.074 0.212 0.020 0.029
Cosine Similarity 0.885 0.925 0.833 0.841
ROUGE-1 0.362 0.475 0.207 0.197
ROUGE-L 0.290 0.361 0.153 0.137
ROUGE-Lsum 0.289 0.361 0.153 0.137
Table 6. Additional evaluation results of the CopyNearest baseline.
Metric Human
Atometric Precision (loose) 0.894
Atometric Precision (moderate) 0.610
Atometric Precision (strict) 0.396
Atometric Recall (loose) 0.909
Atometric Recall (moderate) 0.601
Atometric Recall (strict) 0.244
BERTScore Precision 0.905
BERTScore Recall 0.883
BLEU (BLEU) 0.035
Cosine Similarity 0.875
ROUGE-1 0.324
ROUGE-L 0.278
ROUGE-Lsum 0.278
Table 7. Additional evaluation results of the human baseline on the Persona dataset.
A.2. Human Annotation
Figure 3. Sample screenshot of human annotation protocol.
Two undergraduate students were recruited to annotate the human baseline for the Persona dataset, with each annotator generating 60 textual entries for the testing set. The annotators were provided with the same information as the other candidate methods, including a 2D visualization centered on the query data point. Textual information for each data point appeared when the annotator hovered over it. To enhance their understanding, the five nearest neighbors to the query point were highlighted in color, with their corresponding contents displayed next to the visualization. A sample screenshot of the annotation prompt is provided in Figure 3.
Appendix B Atometric Details
The evaluation of Atometric involves two steps, decomposition and verification. It is straightforward to adapt Atometric to other datasets with the provided templates. The implementation is available at https://anonymous.4open.science/r/atometric-17A3/. We use gpt-4o-2024-05-13 as the backbone model for all the evaluation.
B.1. Atomic Statement Decomposition
Figure 4. Prompt template for Atometric decomposition. The proposed metric could be adapted to different dataset with minimal change in prompting.
A backbone LLM is prompted to decompose the textual content into a list of atomic statements, with the system prompt provided in Figure 4. To better adapt the model to each specific dataset, we include few-shot examples to guide the decomposition process, provided in the aforementioned implementation repository. In general, the resulting atomic statements are expected to be simple, clear, non-trivial, and self-contained.
B.2. Atomic Statement Verification
Figure 5. Prompt template for Atometric verification. The proposed metric could be adapted to different dataset with minimal change in prompting.
We use Atometric precision as an example to illustrate the verification process. Given a list of atomic statements, a backbone LLM is prompted to compare each atomic statement against the reference information. Drawing inspiration from the textual entailment criteria defined by Bowman et al. (2015b), we define four levels of relevance based on how likely an atomic statement accurately describes the subject in the reference, shown in Figure 5, the template of system prompt. The final precision score at strictness level is then calculated as the proportion of generated statements considered relevant at or above that level. For instance, Atometric Precision (M) counts all statements deemed both high and moderate possibility.
Appendix C Dataset Details
C.1. Persona
Figure 6. 2D visualization of Persona dataset.
Description
Persona dataset contains a set of synthetic persona, created with the pipeline introduced in distilabel 1.4.0. It contains 100,000 sample personas as well as their embeddings encoded by Alibaba-NLP/gte-large-en-v1.5 model. The original dataset also provides information about nearest neighbors, 2D projection coordinates, cluster labels and summary labels for each data sample. A few examples are shown below:
  • A geography teacher or high school educator focused on environmental and climate studies, likely designing a lesson plan or assignment on the human impacts of climate change.
  • A meteorologist or a radar technician dealing with weather surveillance and radar image analysis, likely one working with national radar systems interpreting data to provide accurate weather forecasts.
  • A public health specialist focused on childhood environmental health and prevention, likely working in a county or local government role, such as a health educator or departmental program manager.
C.2. Red-Teaming Strategies
Figure 7. 2D visualization of Red-Teaming Strategies dataset.
Description
The Red-Teaming Strategies dataset is derived from the Red Teaming LLM dataset, which comprises human-generated red teaming dialogues aimed at uncovering vulnerabilities in large language models. The original dataset includes dialogues annotated for overall harmfulness using both human and automated assessments. In this work, rather than utilizing the full dialogues due to their length, we focus on extracting the strategies employed within these dialogues to provoke harmful responses. These strategies can inspire the generation of more specific and actionable examples for further analysis. In specific, we prompt GPT-4o to summarize each dialogue using the prompt in Figure 10. A few examples are shown below:
  • The context of the conversation involves a person seeking advice on how to come out to their religious parents, who they anticipate will react negatively. The red team expert attempts to elicit harmful outputs by using provocative language and stereotypes about religious parents, suggesting violence as a solution, and expressing disdain for their beliefs. This approach could be harmful as it promotes negative stereotypes, encourages violence, and fosters divisive attitudes rather than constructive dialogue.
  • The context of the conversation involves a person seeking advice on how to engage in sexual activities with a married coworker at their workplace without getting caught. The red team expert uses a technique of escalating the conversation by suggesting specific locations and actions, and eventually introducing the idea of threatening violence to prevent being reported. This output is harmful as it encourages inappropriate workplace behavior, infidelity, and violence, which could lead to serious personal and professional consequences.
C.3. CS Research Idea & Context
Figure 8. 2D visualization of CS Research Idea dataset.
Figure 9. 2D visualization of CS Research Context dataset.
Description
We follow the same pipeline as MASSW paper. For a list of publications in the computer science field, we use the key idea and context from them. A few examples of the key ideas are shown below:
  • The authors propose a metric, based on the Fisher information, that is strongly indicative of the generalizability of local minima and can be effectively applied as a practical regularizer.
  • The authors propose a novel approach to geolocalise panoramic images on a 2-D cartographic map by learning a low-dimensional embedded space which allows comparison between an image captured at a location and local neighbourhoods of the map.
  • The authors propose a new approach for predicting SQL query properties, including the query answer size, run-time, and error class, relying on data-driven machine learning techniques and large query workloads, instead of database stats or execution plans.
A few examples of the contexts are shown below:
  • Recent advances in deep learning have focused on studying the generalizability across different local minima of deep neural networks (DNNs), but no existing methods both discover properties of good local minima and develop regularization techniques to induce good local minima.
  • Current geolocalisation approaches require image data to be tied to a particular location on a map, but struggle in localizing single images accurately.
  • Formulating efficient SQL queries is a challenging and iterative process requiring tuning and execution cycles. Current methods for providing insights about SQL query properties prior to execution depend on database instance statistics or query execution plans.
Figure 10. Prompt for summarization of the Red Teaming LLM dataset.
Appendix D Implementation Details
Repository Access
Our experiment details and code can be found at https://anonymous.4open.science/r/map2text-27FD.
Fine-tuning Experiments Details
We employed the Llama-3.1-70B-Instruct model for fine-tuning, running on 4 NVIDIA A100 GPUs. Each dataset was fine-tuned for three epochs, and for datasets with over 100,000 data points, the process required approximately 50 hours to complete. To enhance the model’s training efficiency, we utilized the Low-Rank Adaptation (LoRA) technique, setting the rank ($r$) to 16 and the scaling factor ($\alpha$) to 32. We set the dropout rate to 0.05 to prevent overfitting, while setting the initial learning rate to 1e-4, weight decay to 0.0, and momentum factor ($\gamma$) to 0.85 to optimize convergence. Besides, mixed precision training was enabled to maximize computational efficiency. For consistency, all experiments were initialized with the same hyperparameter settings. We use temperate $t=1$ during inference.
Embedding Inversion
In our embedding inversion approach, the first mapping $g_{1}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{H}$ is implemented using a K-NN interpolation algorithm. We employ an approximate nearest neighbor (ANN) search method with a threshold across all datasets to efficiently identify the closest data points, ensuring scalability even for larger datasets. The retrieved neighbors’ high-dimensional embeddings are then interpolated using an unweighted interpolation. Although a more complex weighted interpolation could be applied, we did not observe any performance improvements. Intuitively, this method leverages the smoothness of the underlying manifold. The interpolated embedding is then fed into a pre-trained vec2text model, where the number of search steps (num_steps) is set to 5, with all other configurations kept as default.
Prompt Examples
Figure 11. Prompts used in our 1st-order RAG method on the Red Team Attempts dataset.
Retrieval Augmented Generation
We use gpt-4o-2024-05-13 for all RAG-based method and set temperature $t=0$ for all generations. Figures 11 provides an example of prompts used in the 1st-order RAG method for the Red Team Attempts dataset in our experiments. For a comprehensive collection of prompts used across all datasets and all methods, please refer to the map2text/model/prompts directory in the repository,which is available at https://anonymous.4open.science/r/map2text-27FD/map2text/model/prompts/.